OpenStudy (anonymous):

Help please :c Daniel needs to study a possible link between anti-inflammatory drugs and kidney disease. Help Daniel design a study that would be appropriate. Explain the type of study, the sampling method, the parameter of interest, and the statistic he could use to estimate that parameter. What conclusions would Daniel be able to draw from this study?

OpenStudy (anonymous):

What I came up with was that he basically needs to do a survey rather than an experiment. It seems more logical due to the fact that what if it really did, then now people get kidney disease due to his experiment.

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Preetha @dan815

OpenStudy (anonymous):

Any ideas, Dan?

OpenStudy (dan815):

What are the other types of study, and sampling methods

OpenStudy (dan815):

and parameter of interest--->what age groups to pick from and what sex, background body types and stuff

OpenStudy (anonymous):

Okay, well to make it more legit the parameter of interest should be random as long as they have kidney disease correct?

OpenStudy (anonymous):

It's not like he should stick to 25-30 year old Asians. It should be very open I would imagine.

OpenStudy (dan815):

Yeah I suppose, random sampling method, but parameter of interest maybe an even distribution among the ages, and a random selection in the age groups

OpenStudy (dan815):

To do proper study though

OpenStudy (dan815):

Divide into even age groups, then redivide the age groups by some other characteristic... like weight lets say. And then pick random from this arranged groups

OpenStudy (dan815):

random from each subdivided groups

OpenStudy (dan815):

just random suggestions dont have to follow, im just giving you ideas to organize data.

OpenStudy (anonymous):

Okay, I think I can make something up then. But what about the conclusion, am I suppose to just make something up?

OpenStudy (dan815):

So like the way, he arranges groups lets say... by age or something, he can see what type of age groups are more heavily affected if so... then he can conclude age is afactor and account for it, if not, he can conclude there are other factors and look at other arrangement of data

OpenStudy (anonymous):

Okay, I think I understand. But is it just me who thinks the question isn't completely logical?

OpenStudy (dan815):

It's a wide question, but its fine I guess.

OpenStudy (dan815):

Different possible answers based on your argument.

OpenStudy (anonymous):

Okay. Do you mind helping with another?

OpenStudy (anonymous):

I can post it here or in a new question so I can give you another medal.

OpenStudy (anonymous):

Two gears are connected and are rotating simultaneously. The smaller gear has a radius of 4 inches, and the larger gear has a radius of 7 inches. two circles touching at one point. Larger circle has radius of 7 inches. Smaller circle has radius of 4 inches. Part 1: What is the angle measure, in degrees and rounded to the nearest tenth, through which the larger gear has rotated when the smaller gear has made one complete rotation? Part 2: How many rotations will the smaller gear make during one complete rotation of the larger gear?

OpenStudy (dan815):

okay

OpenStudy (dan815):

Lets look at circles assume the gear spokes are too small to play a factor

OpenStudy (dan815):



OpenStudy (anonymous):

help daniel lol

OpenStudy (dan815):

lets look at 1 full revolution of the small circle, 2*pi*4=8pi, when it has rotated 1 revolution, how much has the big one rotated,

OpenStudy (dan815):

The Big one must have also moved 8pi, however 8pi/(7*2*pi)= 8/14=4/7

OpenStudy (dan815):

You have seen only 4/7 change in rotation

OpenStudy (anonymous):

So while the small gear does a full rotation the big gear does 4/7 of a rotation

OpenStudy (dan815):

We can generalize a basic equation now, for and given radius x and radius y for 2 circles, The angle of rotation change is proportional to x/y for 1 revolution of the change in the circle with radius x

OpenStudy (dan815):

Simply saying... divide the 2 radius to know the change in rotation of the other circle

OpenStudy (anonymous):

So I would divide 7 by 2 is what you're saying?

OpenStudy (anonymous):

Sorry, math is not my strongest subject.

OpenStudy (dan815):

Part 2) let x be rotation of small circles Let y be rotation of big circles We saw that y=4x/7 when y = 1 1=4x/7 x=7/4

OpenStudy (dan815):

Also you are not done part 1.. We saw that it was rotation 4/7 th of the smaller circles rotation, so 4/7 * 360 degrees = the rotation in degrees

OpenStudy (dan815):

Well give all that a good read

OpenStudy (anonymous):

So 205.7 is the rotations in degrees, correct?

OpenStudy (dan815):

Do not worry about the answer, focus on understand the concept, and how you can solve it yourself.

OpenStudy (dan815):

understanding*

OpenStudy (anonymous):

okay.. But I mean either way I have to get the answer.

OpenStudy (dan815):

*flicks hair back*

OpenStudy (dan815):

No rawr the answer is not the goal here.

OpenStudy (anonymous):

.-. I'm just going to guess I was right for the first part. So what you said up there about part two I have to keep solving, correct?

OpenStudy (anonymous):

Okay.

OpenStudy (dan815):

I am trying to help you. You have to be able to solve them on your own. What about doing this question? 2 gears, radius 2 and 5, can you do the same 2 questions in Part 1 and 2 for these 2 gears

OpenStudy (anonymous):

well the rotation fraction of the big circle would be 4/10 right?

OpenStudy (dan815):

yes 2/5 = 4/10

OpenStudy (dan815):

What will be that angle of rotation of the big gear then for 1 full smaller gears rotation

OpenStudy (anonymous):

I would take 4/10 and multiply it by 360 right?

OpenStudy (dan815):

yep good.

OpenStudy (anonymous):

I typed it into a basic calculator so this may be wrong but I got 1.256

OpenStudy (dan815):

1.256 for what?

OpenStudy (anonymous):

Just did it on mathway and got 144.

OpenStudy (dan815):

ok

OpenStudy (anonymous):

So for part 1 of your example my answer would be 144.

OpenStudy (anonymous):

So now we need to do part 2.